OpenStudy (johnny1423):

How do I prove the following properties of partial derivatives? (∂z/∂x)_y=〖[∂x/dz]_y〗^(-1) and (∂x/∂y)_z (∂y/∂z)_x (∂z/∂x)_y=-1

OpenStudy (jango_in_dtown):

High, kindly use the equation mode.

OpenStudy (holsteremission):

Looks like \[\begin{align*} \left(\frac{\partial z}{\partial x}\right)_y&=\left(\left(\frac{\partial x}{\partial z}\right)_y\right)^{-1}&(1)\\[1ex] \left(\frac{\partial x}{\partial y}\right)_z\left(\frac{\partial y}{\partial z}\right)_x\left(\frac{\partial z}{\partial x}\right)_y&=-1&(2) \end{align*}\]though I'm not sure what the subscripts mean. Probably that \(\left(\dfrac{\partial x_i}{\partial x_j}\right)_{x_k}=\dfrac{\partial x_i}{\partial x_j~\partial x_k}\)?

OpenStudy (mhchen):

Can you post a picture of the equation?

OpenStudy (johnny1423):

Sorry for the inconvenient typing, but I don't know how to use the equation mode when typing a question! HolsterEmission these equations are correct. The subscripts represent the variable that is considered a constant in a partial derivative, but I suppose they are not always necessary.

OpenStudy (irishboy123):

assuming \(z = z(x,y)\) and if y = const you can say that \(z = z(x)\) so you have the usual proof that \(y' = \dfrac{1}{x'}\), using decomposition and if it's that limited, the second bit should follow too

OpenStudy (johnny1423):

Still having trouble proving the second property!